# Nonlinear morphoelastic energy based theory for stimuli responsive elastic shells

### v2

@@ 1 Introduction
Extreme deformations are a defining feature of morphing in active and biological slender structures, enabling them to perform their functions. These deformations typically involve both large strains and rotations. In active systems, shape changes are driven by the coupling between elastic responses and non-mechanical stimuli. A notable example is the eversion of the Volvox embryo [Hohn2015], which combines large rotations with active contraction, making it a reference problem for understanding the mechanics of morphing thin structures.
@@
The mechanics of slender shells has been studied extensively, with an important distinction between fluid-like shells, which cannot sustain shear stresses, and solid-like shells, which can. This work focuses on solid-like shells. A widely used constitutive model for passive slender shells is Koiter’s linear theory [Koiter1966], which assumes small strains and moderate rotations. Despite not being rigorously derived as a limit of three-dimensional elasticity (a step later achieved by Steigmann [Steigmann2013]), its popularity stems from the separation of area change and curvature change in the mid-surface. However, Koiter’s model rests on two contrasting assumptions: the Kirchhoff–Love kinematic hypothesis, equivalent to a plane-strain condition, and the mechanical plane-stress condition [Efrati2009, Ciarlet2005]. Later works have sought to resolve these inconsistencies, providing stronger theoretical foundations and incorporating nonlinearities.
@@
Deseri et al. [Desseri2008] rigorously derived a reduced free energy establishing a consistent framework that connects local and non-local contributions and provides a solid basis for modeling biological and elastic membranes. For quadratic energy densities, Wood and Hannah [Wood2019] showed that different strain measures yield different results. More recently, Vitral and Hannah [Vitral2023, Vitral2023b] emphasized the need to include the intrinsic coupling between stretching, bending, and geometry, absent from Koiter’s model but guaranteed when reduction is performed from a fully nonlinear setting [Steigmann2013]. To address the limitations of the classical assumptions, generalized Kirchhoff–Love hypotheses have been proposed [Vitral2023, Ozenda2021, Stumpf1986], allowing more general through-the-thickness responses.
@@
Morphing induced by activity can be incorporated in various ways. When the relaxation of elastic incompatibility is fast compared to the timescale of the stimulus, morphoelasticity [Rodriguez1994, Goriely2017] provides an effective modeling framework, with stimuli introduced through energy functionals [Kumar2023]. Its application to shells began with Efrati et al. [Efrati2009], who extended the linearized theory to non-Euclidean shells. Within this framework, Pezzulla et al. [Pezzulla2018PRL] showed that non-elastic stimuli can act as boundary terms in the variational formulation. Later extensions included large rotations [Haas2021] and fully nonlinear morphoelastic theories derived directly from three dimensions [Sadik2016, Chen2025, Yu2025, Andrini2025].
@@
In this work, starting from a three-dimensional nonlinear morphoelastic theory, we derive a reduced two-dimensional energy functional that governs the morphing of slender, naturally curved objects. We focus on two classical constitutive laws: the incompressible Neo-Hookean material and the compressible Ciarlet–Geymonat model. By retaining geometric and constitutive nonlinearities without additional kinematic or constitutive assumptions, we include the effect of incompatible through-the-thickness stimuli, extending the work of Ozenda and Virga [Ozenda2021] to curved morphoelastic shells. Our formulation is expressed in the language of differential geometry, with the non-elastic stimulus introduced in a general, phenomenological manner.
@@ 2 Non-linear morphoelastic theory
We consider a three-dimensional body $\mathcal{B}_{0}$ embedded in the Euclidean space $\mathcal{E}$, described in terms of a curvilinear material coordinate system $\mathbf{X}=[X^{1},X^{2},X^{3}]$. The position of the body in the reference configuration is given by a mapping $\mathbf{R}(\mathbf{X}):\mathcal{B}_{0}\rightarrow\mathcal{E}$. The associated reference metric is denoted by $\overset{\circ}{\mathrm{G}}_{ij}$, with components defined as $\overset{\circ}{\mathrm{G}}_{ij}=\overset{\circ}{\mathbf{g}}_{i}\cdot\overset{\circ}{\mathbf{g}}_{j}$, where the covariant tangent vectors are $\overset{\circ}{\mathbf{g}}_{i}=\partial\mathbf{R}/\partial X^{i}$. Here and in the following, symbols in bold denote tensorial quantities while symbols in italics denote the component indicated by the subscripts.
The motion of the body is described by a mapping $\chi:\mathcal{B}_{0}\rightarrow\mathcal{E}$, which determines the current configuration $\mathcal{B}$ so that the deformed position vector is given by $\mathbf{r}=\chi(\mathbf{R}(\mathbf{X}))$. The metric associated with the current configuration is denoted by $\mathrm{G}_{ij}$, with components $\mathrm{G}_{ij}=\mathbf{g}_{i}\cdot\mathbf{g}_{j}$, where $\mathbf{g}_{i}=\partial\mathbf{r}/\partial X^{i}$ are the covariant tangent vectors in the deformed configuration. The deformation gradient associated with the motion $\chi$ is defined as $\mathbf{F}=\nabla\chi$, where $\nabla$ denotes the three dimensional gradient with respect to the material coordinates $\mathbf{X}$.
Following the framework of morphoelasticity [Rodriguez1994, Goriely2017], we introduce an intermediate configuration $\bar{\mathcal{B}}$, which may be geometrically incompatible, that is characterized by a prescribed metric $\bar{\mathrm{G}}_{ij}$. This configuration accounts for the presence of a non-elastic stimulus applied to the system. As standard within this framework, we assume a multiplicative decomposition of the deformation gradient tensor as described in fig. 1 and given by
@@
where repeating indices implies summation over $i\in\{1,2,3\}$, and $\bar{\mathbf{g}}^{i}$, $\mathbf{g}^{i}$, and $\overset{\circ}{\mathbf{g}}^{i}$ denote the contravariant basis vectors of the intermediate, current, and reference configurations, respectively. The deformation gradient tensors are defined as
@@
where $\mu^{*}$ is the shear modulus and $p$ is a Lagrange multiplier enforcing the incompressibility constraint, and the compressible Ciarlet–Geymonat material, with energy density [Ozenda2021]
@@
where $\lambda^{*}$ denotes the bulk modulus. We assume that the non-elastic stimulus does not alter the intrinsic material properties (e.g., $\mu^{*}$ and $\lambda^{*}$ remain constant), being consistent with a density-preserving non-elastic stimulus [Lee2023].
@@
where $\delta^{i}_{l}$ is the Kronecker delta and $\bar{\mathrm{G}}^{ij}$ denotes the components of the inverse (contravariant) of the intermediate metric $\bar{\bf G}_{ij}$. The second invariant is written as
@@
$\Psi_{nh}=\int_{\mathcal{B}_{0}}\Phi_{nh}\,J_{m}\,\mathrm{d}X^{1}\mathrm{d}X^{2}\mathrm{d}X^{3},\quad\text{and}\quad\Psi_{cg}=\int_{\mathcal{B}_{0}}\Phi_{cg}\,J_{m}\,\mathrm{d}X^{1}\mathrm{d}X^{2}\mathrm{d}X^{3}\,.$ (11)
@@
Following the consistent finite-strain shell theory derived from 3D nonlinear elasticity as presented by, among others, Ciarlet [Ciarlet2005] and Steigmann [Steigmann2012, Steigmann2013], we first present the kinematics of the shell and describe the geometry of the reference and current configurations.
@@
The formalism employed in our formulation follows the works of Vitral and Hanna [Vitral2023] and Ozenda and Virga [Ozenda2021]. We carefully discuss where our approach diverges from theirs. In particular, we use the description proposed by Ozenda and Virga [Ozenda2021] to relax the plane-strain hypothesis, allowing for changes of the thickness via the introduction of a function $\phi(\eta^{\alpha},Z)$ in the definition of the current configuration as
@@
where $\bf s$} is the position of a point in the current midsurface and $\bf n$} is the unit normal to the current midsurface $\bf s$}. We assume that $\phi\sim O(h_{0})$, so that it can be later expanded as a polynomial function in $Z$, with coefficients determined by minimizing a suitable functional, from either kinematical or energetic arguments. Compared to the work by Ozenda and Virga [Ozenda2021], our problem involves naturally curved shells also subjected to a non-elastic stimulus.
By direct computation, using the ansatz introduced in equation (12), the 3D metric describing the body in the reference configuration takes the simplified form
$\overset{\circ}{\mathrm{G}}_{ij}\mathrm{d}X_{i}\mathrm{d}X_{j}=\overset{\circ}{\mathrm{G}}_{\alpha\beta}\mathrm{d}\eta_{\alpha}\mathrm{d}\eta_{\beta}+Z^{2}\mathrm{d}Z\mathrm{d}Z\,$, (14)
where the metric is decoupled in the in-plane contribution, denoted by $\overset{\circ}{\mathrm{G}}_{\alpha\beta}$, with the Greek indices $\alpha,\beta=1,2$, and the through-the-thickness contribution. The components of the metric associated with those two contributions take the form [Ciarlet2000]
@@
Similarly, by direct computation using the position field in equation (13), the covariant metric for the current configuration can be written as
$\mathrm{G}_{IJ}\mathrm{d}X_{I}\mathrm{d}X_{J}=\mathrm{G}_{\alpha\beta}\mathrm{d}\eta_{\alpha}\mathrm{d}\eta_{\beta}+\phi^{2}\mathrm{d}Z\mathrm{d}Z\,$, (16)
@@
where $\left(\cdot\right)^{\prime}$ indicates differentiation with respect to $Z$ and $\mathrm{a}_{\alpha\beta}$, $\mathrm{b}_{\alpha\beta}$, and $\mathrm{c}_{\alpha\beta}$ are the components of the first, second, and third fundamental forms of the mid-surface in the current configuration, respectively. In deriving these expressions, we make use of two assumptions, highlighted in [Ozenda2021] and justified by the slenderness of the shell, that allow us to simplify the analysis. Indeed, $\mathrm{G}_{\alpha 3},\mathrm{G}_{3\alpha}\ll\mathrm{G}_{33}$ because $\phi(\eta^{\alpha},0)=0$, which implies $\nabla_{\eta}\phi(\eta^{\alpha},0)=0$, and the thickness is small enough to guarantee that $\nabla_{\eta}\phi(\eta^{\alpha},Z)\ll\partial\phi/\partial Z$. Furthermore, since we consider orientation-preserving mappings, we require $\phi(\eta^{\alpha},0)=0$ and $\partial\phi/\partial Z(\eta^{\alpha},0)>0$ that implies that the first coefficient in the (defined later) polynomial expansion of $\phi$ must be positive.
@@
Remark. The coefficient of $Z^{2}$ in equation (3) contains terms that scale with the square of the curvature. Consequently, for consistency, we must retain the contribution of the third fundamental form, which is of the same order as the square of the second fundamental form. In the linearized formulation, as the one based on the derivation proposed by Koiter [Koiter1966], the kinematic ansatz considered implies that the terms of order $O(Z^{2})$ in the metric expansion contribute to the energy only at order $O(h^{5})$. These contributions are therefore neglected in theories that retain only terms of order $O(h)$ and $O(h^{3})$. In contrast, we will show that in our fully nonlinear setting, the order $O(Z^{2})$ terms in the metric expansion contribute at order $O(h^{3})$ to the energy, and thus cannot be ignored.
To model non-elastic stimuli in the mechanics of curved shells, we assume that the metric (not necessarily compatible) associated with the intermediate configuration has the same functional form as in equation (20) so that we write
$\bar{\mathrm{G}}_{IJ}\mathrm{d}X_{I}\mathrm{d}X_{J}=\bar{\mathrm{G}}_{\alpha\beta}\mathrm{d}\eta_{\alpha}\mathrm{d}\eta_{\beta}+\bar{\phi}^{2}\mathrm{d}Z\mathrm{d}Z\,$, (20)
@@
The components of the fundamental forms $\bar{\mathrm{a}}_{\alpha\beta}$, $\bar{\mathrm{b}}_{\alpha\beta}$, and $\bar{c}_{\alpha\beta}$ describe the effect of the non-elastic stimulus in the in-plane direction, while the function $\bar{\phi}(\eta^{\alpha},Z)$ is used to describe its effect across the thickness. We assume that the function $\bar{\phi}$ has all the properties of the function $\phi$: (i) it is of order $O(h)$ and (ii), although the non-elastic stimulus can be of a generic form, it is necessarily associated to an orientation preserving transformation and thus also the first coefficient in the expansion for $\bar{\phi}$ must be positive.
@@
$\bar{\mathrm{G}}^{\alpha\beta}=\bar{\mathrm{a}}^{\alpha\beta}+2\bar{\phi}\bar{\mathrm{a}}^{\alpha\gamma}\bar{\mathrm{b}}_{\gamma\delta}\bar{\mathrm{a}}^{\delta\beta}+3\bar{\phi}^{2}\bar{\mathrm{a}}^{\alpha\gamma}\bar{c}_{\gamma\delta}\bar{\mathrm{a}}^{\delta\beta}+O(\bar{\phi}^{3}),\qquad\mathrm{G}^{33}=\left(\bar{\phi}^{\prime}\right)^{-2}.$ (22)
Remark. In the approach used in Yu and Chen [Yu2025] to derive a two-dimensional model for morphoelastic shells, or in Moulton et al. [Moulton2020] to derive a one-dimensional model for morphoelastic rods, the authors include the non-elastic stimulus by prescribing the deformation gradient $\bar{\bf F}$. Then, by performing a suitable asymptotic expansion of the fields in the slenderness and with respect to the mid-surface or the centerline, the reduced energy can be derived by minimizing a problem on the cross-sections [Lestringant2020]. Our approach introduces the non-elastic stimulus by assuming a particular form of the incompatible metrics as shown in equation (20). This approach translates into the possibility to independently prescribe three quantities: the first and second fundamental forms $\bar{\mathrm{a}}_{\alpha\beta}$ and $\bar{\mathrm{b}}_{\alpha\beta}$, associated with the in-plane contribution of the stimulus, and the function $\bar{\phi}$, associated with the through-the-thickness contribution of the stimulus. The non-elastic stimuli are often difficult to model formally because of the complexity of the physics that regulates their origin. We thus decide to consider a phenomenological approach that, although relying on a specific ansatz for the incompatible metrics as in equation (76), gives the possibility to explore a large variety of non-elastic stimuli. The two formalisms are equivalent and they are connected via the relation
$\bar{\mathbf{C}}=\bar{\mathrm{G}}_{ij}\left(\overset{\circ}{\mathbf{g}}_{i}\otimes\overset{\circ}{\mathbf{g}}^{j}\right),$ (23)
where one can see that it is possible to either prescribe $\bar{\mathbf{C}}$ (or $\bar{\bf F}$) or the metric $\bar{\mathrm{G}}_{ij}$. For example, in the case of an isotropic three dimensional dilation of factor $\xi$, we can write
$\bar{\mathbf{C}}=\xi^{2}\overset{\circ}{\mathrm{G}}_{ij}\left(\overset{\circ}{\mathbf{g}}_{i}\otimes\overset{\circ}{\mathbf{g}}^{j}\right)$ (24)
that directly implies
@@
Given the form in equation (76), it can be directly derived that
@@
$\left(\frac{\partial\bar{\phi}}{\partial Z}\right)^{2}=\xi^{2}$ (27)
@@
(and naturally $\bar{\mathrm{c}}_{\alpha\beta}=\overset{\circ}{\mathrm{c}}_{\alpha\beta}$). Incidentally, we also recover that a uniform dilation, when applied to naturally curved systems, implies the emergence of both stretching of the mid-surface and change of its curvature [Vitral2023].

### v3

@@ 1 Introduction
Extreme deformations are a defining feature of morphing in active and biological slender structures, enabling them to perform their functions. These deformations typically involve both large strains and rotations. In active systems, shape changes are driven by the coupling between elastic responses and non-mechanical stimuli. A notable example is the eversion of the Volvox embryo [13], which combines large rotations with active contraction, making it a reference problem for understanding the mechanics of morphing thin structures.
The mechanics of thin shells has been studied extensively, with an important distinction between fluid-like shells and solid-like shells. The former are characterized by the absence of energetic cost associated with the in-plane rearrangement of the material points; thus, their energy does not depend on in-plane shear deformations, although this does not imply that shear stresses cannot be sustained [7, 29]. On the contrary, the energy associated with the latter depends on the full in-plane metric and not only on its determinant.
@@
This work focuses on solid-like shells. A widely used constitutive model for passive slender shells is Koiter’s shell theory [17, 18], which assumes small strains while retaining geometric nonlinearities associated with moderate rotations. Despite not being rigorously derived as a limit of three-dimensional elasticity (a step later achieved by Steigmann [31]), its popularity stems from the resulting separation between area change and curvature change of the mid-surface in the energy functional proposed. As discussed by Ciarlet [4], Koiter’s model rests on two main assumptions: the Kirchhoff–Love kinematic hypothesis, geometrical in nature, and the plane-stress hypothesis, mechanical in nature. Later works have provided stronger theoretical foundations for these assumptions [15, 16].
@@
Deseri et al. [5] rigorously derived a reduced free energy establishing a consistent framework that connects local and non-local contributions and provides a solid basis for modeling biological and elastic membranes. For quadratic energy densities, Wood and Hanna [36] showed that different strain measures yield different results. More recently, Vitral and Hanna [34, 33] emphasized the presence of the intrinsic coupling between stretching, bending, and geometry, absent in Koiter’s model but guaranteed when reduction is performed from a fully nonlinear setting [31]. To address the limitations of the classical assumptions, generalized Kirchhoff–Love hypotheses have been proposed [34, 24, 32], allowing more general through-the-thickness responses.
@@
Morphing induced by activity can be incorporated in various ways. When the relaxation of elastic incompatibility is fast compared to the timescale of the stimulus, morphoelasticity [27, 10] provides an effective modeling framework, with stimuli introduced through energy functionals [19]. Its application to shells began with Efrati et al. [8], who extended the linearized theory to non-Euclidean shells. Within this framework, Pezzulla et al. [25] showed that non-elastic stimuli can act as boundary terms in the variational formulation. Later extensions included large rotations [11] and fully nonlinear morphoelastic theories derived directly from three dimensions [28, 2, 37, 1].
@@
In this work, starting from a three-dimensional nonlinear morphoelastic theory, we derive a reduced two-dimensional energy functional that governs the morphing of slender, naturally curved objects. We focus on two classical constitutive laws: the incompressible Neo-Hookean material and the compressible Ciarlet–Geymonat model. By retaining geometric and constitutive nonlinearities without additional kinematic or constitutive assumptions, we include the effect of incompatible through-the-thickness stimuli, extending the work of Ozenda and Virga [24] to curved morphoelastic shells. Our formulation is expressed in the language of differential geometry, with the non-elastic stimulus introduced in a general, phenomenological manner.
@@ 2 Non-linear morphoelastic theory
We consider a three-dimensional body $\mathcal{B}_{0}$ embedded in the Euclidean space $\mathcal{E}$, described in terms of a curvilinear material coordinate system $X^{1},X^{2},X^{3}]$. The position of the body in the reference configuration is given by a mapping $\mathbf{R}(X^{i}):\mathcal{B}_{0}\rightarrow\mathcal{E}$ with $i\in\{1,2,3\}$. The associated reference metric is denoted by $\overset{\circ}{\mathrm{G}}_{ij}$, with components defined as $\overset{\circ}{\mathrm{G}}_{ij}=\overset{\circ}{\mathbf{g}}_{i}\cdot\overset{\circ}{\mathbf{g}}_{j}$, where the covariant tangent vectors are $\overset{\circ}{\mathbf{g}}_{i}=\partial\mathbf{R}/\partial X^{i}$. Here and in the following, symbols in bold denote tensorial quantities while symbols in italics denote the component indicated by the subscripts.
The motion of the body is described by a mapping $\chi:\mathcal{B}_{0}\rightarrow\mathcal{E}$, which determines the current configuration $\mathcal{B}$ so that the deformed position vector is given by $\mathbf{r}=\chi(\mathbf{R}((X^{i})))$. The metric associated with the current configuration is denoted by $\mathrm{G}_{ij}$, with components $\mathrm{G}_{ij}=\mathbf{g}_{i}\cdot\mathbf{g}_{j}$, where $\mathbf{g}_{i}=\partial\mathbf{r}/\partial X^{i}$ are the covariant tangent vectors in the deformed configuration. The deformation gradient associated with the motion $\chi$ is defined as $\mathbf{F}=\nabla\chi$, where $\nabla$ denotes the three dimensional gradient with respect to the material coordinates $X^{i}$.
Following the framework of morphoelasticity [27, 10], we introduce an intermediate configuration $\bar{\mathcal{B}}$, which may be geometrically incompatible, that is characterized by a prescribed metric $\bar{\mathrm{G}}_{ij}$. This configuration accounts for the presence of a non-elastic stimulus applied to the system. As standard within this framework, we assume a multiplicative decomposition of the deformation gradient tensor as described in fig. 1 and given by
@@
where repeating indices implies summation over $i$, and $\bar{\mathbf{g}}^{i}$, $\mathbf{g}^{i}$, and $\overset{\circ}{\mathbf{g}}^{i}$ denote the contravariant basis vectors of the intermediate, current, and reference configurations, respectively. We point out that, although the intermediate configuration is not associated with an embedding, it nonetheless describes a locally differentiable manifold so it is still possible to define tangent vectors. The deformation gradient tensors are defined as
@@
where $\mu^{*}$ is the shear modulus and $p$ is a Lagrange multiplier enforcing the incompressibility constraint, and the compressible Ciarlet–Geymonat material, with energy density [24]
@@
where $\lambda^{*}$ denotes the bulk modulus. We assume that the non-elastic stimulus does not alter the intrinsic material properties (e.g., $\mu^{*}$ and $\lambda^{*}$ remain constant), being consistent with a density-preserving non-elastic stimulus [21].
@@
where $\delta^{i}_{l}$ is the Kronecker delta and $\bar{\mathrm{G}}^{ij}$ denotes the components of the inverse (contravariant) of the intermediate metric $\bar{\bf G}_{ij}$. The intermediate metric describes a locally stress-free geometry. In relating the intermediate metrics and the tangent vectors in the intermediate configuration we use $\bar{\mathrm{G}}_{ij}=\bar{\mathbf{g}}_{i}\cdot\bar{\mathbf{g}}_{j}$. The second invariant is written as
@@
$\Psi_{nh}=\int_{\mathcal{B}_{0}}\Phi_{nh}\,J_{m}\sqrt{\det(\overset{\circ}{\mathrm{G}}_{ij})}\,\mathrm{d}X^{1}\mathrm{d}X^{2}\mathrm{d}X^{3},$ (11)
$\Psi_{cg}=\int_{\mathcal{B}_{0}}\Phi_{cg}\,J_{m}\sqrt{\det(\overset{\circ}{\mathrm{G}}_{ij})}\,\mathrm{d}X^{1}\mathrm{d}X^{2}\mathrm{d}X^{3}\,.$
@@
Following the consistent finite-strain shell theory derived from 3D nonlinear elasticity as presented by, among others, Ciarlet [4] and Steigmann [30, 31], we first present the kinematics of the shell and describe the geometry of the reference and current configurations.
@@
The formalism employed in our formulation follows the works of Vitral and Hanna [34] and Ozenda and Virga [24]. We carefully discuss where our approach diverges from theirs. In particular, we use the description proposed by Ozenda and Virga [24] to relax the plane-strain hypothesis, allowing for changes of the thickness via the introduction of a function $\phi(\eta^{\alpha},Z)$ in the definition of the current configuration as
@@
where $\bf s$} is the position of a point in the current midsurface and $\bf n$} is the unit normal to the current midsurface $\bf s$}. We assume that $\phi\sim O(h_{0})$, so that it can be later expanded as a polynomial function in $Z$, with coefficients determined by minimizing a suitable functional, from either kinematical or energetic arguments. Compared to the work by Ozenda and Virga [24], our problem involves naturally curved shells also subjected to a non-elastic stimulus. The use of the kinematic assumption of Kirchhoff-Love type is justified by our interest in considering the mechanics of thin isotropic shells in bending- and membrane-dominated regimes, for which this kinematic relation can be formally derived via asymptotic expansion [3, 4].
By direct computation, using the ansatz introduced in eq.  (12), the 3D metric describing the body in the reference configuration takes the simplified form
$\overset{\circ}{\mathrm{G}}_{ij}\mathrm{d}X^{i}\mathrm{d}X^{j}=\overset{\circ}{\mathrm{G}}_{\alpha\beta}\mathrm{d}\eta^{\alpha}\mathrm{d}\eta^{\beta}+Z^{2}\mathrm{d}Z\mathrm{d}Z\,$, (14)
where the metric is decoupled in the in-plane contribution, denoted by $\overset{\circ}{\mathrm{G}}_{\alpha\beta}$, with the Greek indices $\alpha,\beta=1,2$, and the through-the-thickness contribution. The components of the metric associated with those two contributions take the form [3]
$\overset{\circ}{\mathrm{G}}_{\alpha\beta}=\overset{\circ}{\mathrm{a}}_{\alpha\beta}(\eta^{\alpha})-2Z\overset{\circ}{\mathrm{b}}_{\alpha\beta}(\eta^{\alpha})+Z^{2}\overset{\circ}{\mathrm{c}}_{\alpha\beta}(\eta^{\alpha})\,,\qquad\overset{\circ}{\mathrm{G}}_{33}=1\,$, (15)
@@
Similarly, by direct computation using the position field in eq.  (13), the covariant metric for the current configuration can be written as
$\mathrm{G}_{ij}\mathrm{d}X^{i}\mathrm{d}X^{j}=\mathrm{G}_{\alpha\beta}\mathrm{d}\eta^{\alpha}\mathrm{d}\eta^{\beta}+\phi^{2}\mathrm{d}Z\mathrm{d}Z\,$, (16)
@@
where $\left(\cdot\right)^{\prime}$ indicates differentiation with respect to $Z$ and $\mathrm{a}_{\alpha\beta}$, $\mathrm{b}_{\alpha\beta}$, and $\mathrm{c}_{\alpha\beta}$ are the components of the first, second, and third fundamental forms of the mid-surface in the current configuration, respectively. In deriving these expressions, we make use of two assumptions, highlighted in [24] and justified by the slenderness of the shell, that allow us to simplify the analysis. Indeed, $\mathrm{G}_{\alpha 3},\mathrm{G}_{3\alpha}\ll\mathrm{G}_{33}$ because $\phi(\eta^{\alpha},0)=0$, which implies $\nabla_{\eta}\phi(\eta^{\alpha},0)=0$, and the thickness is small enough to guarantee that $\nabla_{\eta}\phi(\eta^{\alpha},Z)\ll\partial\phi/\partial Z$. Furthermore, since we consider orientation-preserving mappings, we require $\phi(\eta^{\alpha},0)=0$ and $\partial\phi/\partial Z(\eta^{\alpha},0)>0$ that implies that the first coefficient in the (defined later) polynomial expansion of $\phi$ must be positive.
@@
Remark. The coefficient of $Z^{2}$ in eq.  (3) contains terms that scale with the square of the curvature. Consequently, for consistency, we must retain the contribution of the third fundamental form, which is of the same order as the square of the second fundamental form. In the linearized formulation, as the one based on the derivation proposed by Koiter [18], the kinematic ansatz considered implies that the terms of order $O(Z^{2})$ in the metric expansion contribute to the energy only at order $O(h^{5})$. These contributions are therefore neglected in theories that retain only terms of order $O(h)$ and $O(h^{3})$. In contrast, we will show that in our fully nonlinear setting, the order $O(Z^{2})$ terms in the metric expansion contribute at order $O(h^{3})$ to the energy, and thus cannot be ignored.
To model non-elastic stimuli in the mechanics of curved shells, we assume that the metric (not necessarily compatible) associated with the intermediate configuration has the same form as in eq.  (16) so that we write
$\bar{\mathrm{G}}_{IJ}\mathrm{d}X^{I}\mathrm{d}X^{J}=\bar{\mathrm{G}}_{\alpha\beta}\mathrm{d}\eta^{\alpha}\mathrm{d}\eta^{\beta}+\bar{\phi}^{2}\mathrm{d}Z\mathrm{d}Z\,$, (20)
@@
In the intermediate configuration, the components of the target forms $\bar{\mathrm{a}}_{\alpha\beta}$, $\bar{\mathrm{b}}_{\alpha\beta}$, and $\bar{c}_{\alpha\beta}$ describe the effect of the in-plane non-elastic stimulus, while the function $\bar{\phi}(\eta^{\alpha},Z)$ is used to describe its effect across the thickness. It is important to note that the target forms do not have the same meaning as the fundamental forms introduced when we presented the expansion of the metrics in the reference and current configurations: they rather must be seen as coefficients of an appropriate asymptotic expansion of the intermediate metric and, as a consequence, $\bar{c}_{\alpha\beta}$ is not necessarily derived from $\bar{\mathrm{a}}_{\alpha\beta}$ and $\bar{\mathrm{b}}_{\alpha\beta}$, contrarily to the third fundamental form being dependent on the first and the second ones in the reference and current configurations that are both associated with an embedding. Later, to simplify the analysis, we introduce constraints on $\bar{c}_{\alpha\beta}$ and we leave the analysis of the general form to future work. Furthermore, we assume that the function $\bar{\phi}$ has all the properties of the function $\phi$: (i) it is of order $O(h)$ and (ii), although the non-elastic stimulus can be of a generic form, it is necessarily associated to an orientation preserving transformation and thus also the first coefficient in the expansion for $\bar{\phi}$ must be positive.
@@
$\bar{\mathrm{G}}^{\alpha\beta}=\bar{\mathrm{a}}^{\alpha\beta}+2\bar{\phi}\bar{\mathrm{a}}^{\alpha\gamma}\bar{\mathrm{b}}_{\gamma\delta}\bar{\mathrm{a}}^{\delta\beta}+3\bar{\phi}^{2}\bar{\mathrm{a}}^{\alpha\gamma}\bar{c}_{\gamma\delta}\bar{\mathrm{a}}^{\delta\beta}+O(\bar{\phi}^{3})\,,\qquad\bar{\mathrm{G}}^{33}=\left(\bar{\phi}^{\prime}\right)^{-2}\,.$ (22)
The form presented in eq. (22) introduces a constraint on $\bar{c}_{\alpha\beta}$ because it is derived by forcing $\bar{c}_{\alpha\beta}=\bar{\mathrm{b}}_{\alpha\gamma}\bar{\mathrm{a}}^{\gamma\delta}\bar{\mathrm{b}}_{\delta\beta}$ and, from now on we force this requirement. This does not affect the characteristic of the intermediate configuration of being globally incompatible because, setting $\bar{\mathrm{a}}_{\alpha\beta}$ and $\bar{\mathrm{b}}_{\alpha\beta}$ arbitrarily, we still guarantee to impose a target metric that is not necessarily associated with an embedding.
Remark. In the approach used in Yu and Chen [37] to derive a two-dimensional model for morphoelastic shells, or in Moulton et al. [23] to derive a one-dimensional model for morphoelastic rods, the authors include the non-elastic stimulus by prescribing the deformation gradient $\bar{\bf F}$. Then, by performing a suitable asymptotic expansion of the fields in the slenderness and with respect to the mid-surface or the centerline, the reduced energy can be derived by minimizing a problem on the cross-sections [22]. Our approach introduces the non-elastic stimulus by assuming a particular form of the incompatible metrics as shown in eq.  (20). This approach translates into the possibility to independently prescribe three quantities: the first and second fundamental forms $\bar{\mathrm{a}}_{\alpha\beta}$ and $\bar{\mathrm{b}}_{\alpha\beta}$, associated with the in-plane contribution of the stimulus, and the function $\bar{\phi}$, associated with the through-the-thickness contribution of the stimulus. The non-elastic stimuli are often difficult to model formally because of the complexity of the physics that regulates their origin. We thus decide to consider a phenomenological approach that, although relying on a specific ansatz for the incompatible metrics as in eq.  (21), gives the possibility to explore a large variety of non-elastic stimuli. The two formalisms are equivalent and they are connected via the relation
$\bar{\mathbf{C}}=\bar{\mathrm{G}}_{ij}\left(\overset{\circ}{\mathbf{g}}^{{i}}\otimes\overset{\circ}{\mathbf{g}}^{j}\right)\,$, (23)
where one can see that it is possible to either prescribe $\bar{\mathbf{C}}$ (or $\bar{\bf F}$) or the metric $\bar{\mathrm{G}}_{ij}$ and derive the other. We illustrate this by looking at a simple non-mechanical stimulus of an isotropic three dimensional dilation of factor $\xi$; we can write
$\bar{\mathbf{C}}=\xi^{2}\overset{\circ}{\mathrm{G}}_{ij}\left(\overset{\circ}{\mathbf{g}}^{{i}}\otimes\overset{\circ}{\mathbf{g}}^{j}\right)\,$, (24)
which directly implies
@@
Given the form in eq.  (21), it can be directly derived that
@@
$\left(\frac{\partial\bar{\phi}}{\partial Z}\right)^{2}=\xi^{2}\,$, (27)
@@
(and naturally $\bar{\mathrm{c}}_{\alpha\beta}=\overset{\circ}{\mathrm{c}}_{\alpha\beta}$). We note that this stimulus, chosen here for its simplicity, does not describe an incompatible intermediate configuration but rather gives rise to a realizable manifold in the Euclidean space without the need of any compatibility to be restored by the elastic contribution, provided that boundary conditions allow it. Inspection of eq. (28) shows that an isotropic dilation in curved shells, $\overset{\circ}{\mathrm{b}}_{\alpha\beta}\neq 0$, is not just pure stretching because you must include a constant through-the-thickness stretch that causes the curvature to also change [34].